# PolyominoIdeals: a package for Macaulay2 to work with the inner $2$-minor ideals of collections of cells

### v1

@@ 1. Introduction
An unitary square in $\mathbb{Z}^{2}$, whose vertices have integer coordinates, is said to be a cell. Considering collections of cells in the plane, it is possible to create many geometric objects having a different shapes. A particular class of collections of cells consists of polyominoes. A polyomino is a finite collection of unitary squares joined edge by edge. This term was coined by S. W. Golomb in 1953 and such configurations have been studied mainly in Combinatorial Mathematics, for instance in some tiling problems of the plane. In 2012 A.A. Qureshi (see [27]) defined a polyomino ideal attached to a polyomino $\mathcal{P}$, as the ideal generated by all inner 2-minors of $\mathcal{P}$ in the polynomial ring $S_{\mathcal{P}}$, that is the polynomial ring over a field $K$ in the variables $x_{v}$, for all $v$ vertices of cells in $\mathcal{P}$. Such an ideal is called the polyomino ideal of $\mathcal{P}$ and it is denoted by $I_{\mathcal{P}}$. For more details see [27]. The study of the main algebraic properties of the quotient ring $K[\mathcal{P}]=S_{\mathcal{P}}/I_{\mathcal{P}}$, depending on the shape of $\mathcal{P}$, has become an exciting line of research. Many authors have focused their attention about that, see [1], [2], [3], [4], [5], [6], [7], [8], [9], [13], [15], [16], [17], [20], [21], [23], [24], [25], [28], [29], [30], [31].
@@
In this paper, we provide a description of the package PolyominoIdeals for the computer algebra software Macaulay2. The purpose of this package is to allow the definition and manipulation of the binomial ideals attached to collections of cells, inside the computer algebra software. More precisely, encoding a fixed collection of cells by a list of lists containing the diagonal corners of each cell, the package provides three functions. The function polyoIdeal allows to define the inner $2$-minor ideal of $\mathcal{P}$. For this function, we also give three options. The option RingChoice allows one to choose between two rings having two different monomial orders: one is the lexicographic order induced by the natural partial order on $V(\mathcal{P})$; the other one is the monomial order $<$ given in [18] for bipartite graphs, that can be induced in a polyomino as shown in [27], where one also can deduce that, for a convex polyomino $\mathcal{P}$, the generators of the ideal $I_{\mathcal{P}}$ form the reduced Gröbner basis with respect to $<$. For the other two options, the first one is Field which allows changing the base field $K$ in the base polynomial ring of $I_{\mathcal{P}}$, and the second is TermOder, which allows replacing the lexicographic order by other term orders when RingChoice is defined in the first way as described above. For the other functions, we have polyoToric, which allows generating the toric ideal defined in [24]. The third function polyoMatrix gives the matrix attached to $\mathcal{P}$, that is the matrix whose set of all $2$-minors consists of the set of generators of the ideal $I_{\mathcal{P}}$. The paper is arranged as follows. In Section 2 we introduce and explain in detail the main mathematical definitions for collections of cells and the associated ideals, needed for what we describe in this work. In Section 3 we describe the functions and the related options and we show several examples of them. In Section 4 we explain a useful trick that allows getting the list encoding the collection of cells using GeoGebra.
@@ 3. Functions in the package PolyoIdeals
In this section we describe the functions included in the package PolyominoIdeals.m2 developed for Macaulay2 ([12]). First of all, consider that a collection of cells $\mathcal{P}$ is encoded, for the package, with a list of lists, where each list represents a cell of the collection and contains two lists representing the diagonal corners of the cell, the first for the lower left corner, the second for the upper right corner. For instance, the collection of cells in Figure 2 is encoded with the list Q = {{{1, 1}, {2, 2}}, {{2, 1}, {3, 2}}, {{3, 1}, {4, 2}}, {{2, 2}, {3, 3}}, {{3, 2}, {4, 3}}, {{2, 3}, {3, 4}}}.
@@
Figure 2. A polyomino.
@@ 3.6. An example of computing other properties of the polyomino ideal
4. A trick with the help of GeoGebra
In order to make tests for some big polyominoes or collections of cells, it could be not so easy to provide to Macaulay2 the list of the diagonal corners of each cell that encodes our object. In this section, we give a trick to obtain the encoding of a desired collection of cells, after drawing it in GeoGebra. GeoGebra is a dynamic mathematics software for all levels of education that brings together geometry, algebra, spreadsheets, graphing, statistics, and calculus in one engine. Its usage is very intuitive, in particular in the Graphics View it is possible to draw points, vectors, segments, lines, polygons, and many other things. The Graphics View Toolbar provides a wide range of Tools that allow you to create the graphical representations of objects directly in the Graphics View. GeoGebra allows to export its materials in several formats, including SVG, Animated GIF, Windows Metafile, PNG, PDF, and EPS, and can also generate code for use in files. GeoGebra is available on multiple platforms, with apps for desktops (Windows, macOS and Linux), tablets (Android, iPad and Windows) and web. For many other insights see [10]. In the following explanation, we use the web app in the English(UK) language.
First of all, we explain how a collection of cells must be drawn in GeoGebra. In the Graphics View, by default, the coordinate axes and the grid appear. In the Graphics View Toolbar select the tool Regular polygon. This tool works in the following way: if one selects two points $A$ and $B$ in the plane and specifies the number $n$ of vertices in the input field of the appearing dialog window, then a regular polygon with $n$ vertices including points $A$ and $B$ is provided. We use this tool to draw each cell of the collection we want to study, as a regular polygon of $4$ vertices. For our purpose, each cell must be drawn by selecting first the lower left corner, say $A$, and then the lower right corner $A+(1,0)$, using only integer points (see Figure 5).
Figure 5.
In order to continue, we provide here the code of the function “GeoPoly” that we need to import in Macaualy2. If one uses a different language in GeoGebra, the code of the function must be changed opportunely.
    GeoPoly=(n)->(
    out={};
    for i from 1 to n do (
        out=join(out,{”{Vertex(poly”|i|”,1),Vertex(poly”|i|”,3)}”});
    ”vertices” << out << endl << close;
    return out;
Now suppose the collection of cells drawn in GeoGebra has $n$ cells. In the command line of Macaulay2, type the command “GeoPoly($n$)”. This function creates the text file “vertices.txt” in the local folder where Macaulay2 is launched. So, open the file “vertices.txt” and copy and paste its content in the Input Bar of the GeoGebra window (to activate it, go to the options up-right) where the collection of cells is drawn, and type Enter (see Figure 6).
Figure 6.
In this way a list is generated, where each element is the list of the diagonal corners of each cell. Let $l1$ be the name of this list and, in the Input bar, type the command “Text(l1)” (see Figure 7).
Figure 7.
Observe that, at this point, we can obtain the input data in Macaulay2 of the drawn collection of cells by the text appearing in the Graphic view. We only need to change parentheses with braces.
A possible way to obtain the character string of the text in the Graphic view, is to export it as PGF/TikZ (see Figure 8).
Figure 8.
In the exported file (that is a .txt file), one can find the desired character string in a particular line (see Figure 9 for the example explained in the previous pictures).
Figure 9.
Finally, copy the character string, open Macaulay2 and paste it into the command line creating a variable P. That is, in the command line of Macaulay2 you have something like P={{(3, 2), (4, 3)}, {(4, 2), (5, 3)}, {(4, 3), (5, 4)}} (considering the polyomino in the previous pictures). In order to replace parentheses in braces, and so obtain the right encoding of the desired collection of cells in a variable Q, it suffices to type the following commands:
Macaulay2, version 1.20
    with packages: ConwayPolynomials, Elimination, IntegralClosure,
    PrimaryDecomposition, ReesAlgebra, Saturation, TangentCone
i1 : P={{(3, 2), (4, 3)}, {(4, 2), (5, 3)}, {(4, 3), (5, 4)}};
i2 : Q={};
for i from 0 to #P-1 do(
        Q=join(Q,{{toList(P#i#0),toList(P#i#1)}});
i4 : Q
o4 = {{{3, 2}, {4, 3}}, {{4, 2}, {5, 3}}, {{4, 3}, {5, 4}}}
o4 : List
The list Q encodes the desired collection of cells drawn in the Graphic view of GeoGebra, and from it, we are able to use the functions of the package PolyominoIdeals.m2.

### v2

@@ 1. Introduction
An unitary square in $\mathbb{Z}^{2}$, whose vertices have integer coordinates, is said to be a cell. Considering collections of cells in the plane, it is possible to create many geometric objects having a different shapes. A particular class of collections of cells consists of polyominoes. A polyomino is a finite collection of unitary squares joined edge by edge. This term was coined by S. W. Golomb in 1953 and such configurations have been studied mainly in Combinatorial Mathematics, for instance in some tiling problems of the plane. In 2012 A.A. Qureshi (see [20]) defined a polyomino ideal attached to a polyomino $\mathcal{P}$, as the ideal generated by all inner 2-minors of $\mathcal{P}$ in the polynomial ring $S_{\mathcal{P}}$, that is the polynomial ring over a field $K$ in the variables $x_{v}$, for all $v$ vertices of cells in $\mathcal{P}$. Such an ideal is called the polyomino ideal of $\mathcal{P}$ and it is denoted by $I_{\mathcal{P}}$. For more details see [20]. The study of the main algebraic properties of the quotient ring $K[\mathcal{P}]=S_{\mathcal{P}}/I_{\mathcal{P}}$, depending on the shape of $\mathcal{P}$, has become an exciting line of research. Many authors have focused their attention about that, see [2], [3], [4], [5], [6], [7], [8], [12], [13], [14], [17], [18], [19], [21], [22], [23], [24].
@@
In this paper, we provide a description of the package PolyominoIdeals for the computer algebra software Macaulay2. The purpose of this package is to allow the definition and manipulation of the binomial ideals attached to collections of cells, inside the computer algebra software. More precisely, encoding a fixed collection of cells by a list of lists containing the diagonal corners of each cell, the package provides three functions. The function polyoIdeal allows to define the inner $2$-minor ideal of $\mathcal{P}$. For this function, we also give three options. The option RingChoice allows one to choose between two rings having two different monomial orders: one is the lexicographic order induced by the natural partial order on $V(\mathcal{P})$; the other one is the monomial order $<$ given in [15] for bipartite graphs, that can be induced in a polyomino as shown in [20], where one also can deduce that, for a convex polyomino $\mathcal{P}$, the generators of the ideal $I_{\mathcal{P}}$ form the reduced Gröbner basis with respect to $<$. For the other two options, the first one is Field which allows changing the base field $K$ in the base polynomial ring of $I_{\mathcal{P}}$, and the second is TermOder, which allows replacing the lexicographic order by other term orders when RingChoice is defined in the first way as described above. For the other functions, we have polyoToric, which allows generating the toric ideal defined in [18]. The third function polyoMatrix gives the matrix attached to $\mathcal{P}$, that is the matrix whose set of all $2$-minors consists of the set of generators of the ideal $I_{\mathcal{P}}$. The paper is arranged as follows. In Section 2 we introduce and explain in detail the main mathematical definitions for collections of cells and the associated ideals, needed for what we describe in this work. In Section 3 we describe the functions and the related options and we show several examples of them.
In order to make tests for some big polyominoes or collections of cells, it could be not so easy to provide to Macaulay2 the list of the diagonal corners of each cell that encodes our object. Refer to the tool provided in
https://rizwan.phd/polyo/
to draw a collection of cells and to get the desired input data for Macaulay2, that is, the list encoding the pictured collection of cells. See also Section 3.
@@ 3. Functions in the package PolyoIdeals
In this section we describe the functions included in the package PolyominoIdeals.m2 developed for Macaulay2 ([10]). First of all, consider that a collection of cells $\mathcal{P}$ is encoded, for the package, with a list of lists, where each list represents a cell of the collection and contains two lists representing the diagonal corners of the cell, the first for the lower left corner, the second for the upper right corner. Refer to the tool provided in
https://rizwan.phd/polyo/
to draw a collection of cells and to get the desired input data for Macaulay2, that is, the list encoding the pictured collection of cells.
For instance, the collection of cells in Figure 2 is encoded with the list
Q = {{{1, 1}, {2, 2}}, {{2, 1}, {3, 2}}, {{3, 1}, {4, 2}}, {{2, 2}, {3, 3}}, {{3, 2}, {4, 3}}, {{2, 3}, {3, 4}}}
that one can easily obtained drawing the polyomino by clicking on the cells in the grid.
@@
Figure 2. A polyomino and its related list Q.